# Retained-spin micropolar hydrodynamics from the Boltzmann–Curtiss equation: a generalized Chapman–Enskog construction

### 31 Mar 2026 v1

@@ IV.3 Linearized first-order equation
where the source $\mathcal{S}$ collects the $O(\varepsilon)$ streaming contributions together with the matched-order residual axial relaxation term described above. The standard density, temperature, and heat-flux sectors follow the classical Chapman–Enskog pattern [2, 22]. Here we focus on the sectors relevant to the stress and couple stress.
@@ IV.4 Useful derivatives of the local Maxwellian
The last term $\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}$ is written separately because the axial mismatch sector is not generated by the symmetric one-particle kinetic stress alone. In the matched ordering described above, this symbol also absorbs the $O(\varepsilon)$ residual collisional production of the retained-spin quasi-equilibrium manifold. The axial sector therefore belongs to the intrinsic/collisional transfer channel and will be treated formally in Sec. VI.
@@ V.2 Total stress at first order
Since the kernel $c_{i}c_{j}$ is symmetric, only the symmetric sectors in Eq. (81) contribute to $\sigma_{ij}^{(k,1)}$. Rotational invariance then forces the result to be of the form
@@
The antisymmetric channel belongs to the intrinsic/collisional transfer part of the total stress. Introduce the axial mismatch tensor
@@
Isotropy and parity imply that the first-order antisymmetric stress must be proportional to $A_{ij}$, so we define
$\sigma_{[ij]}^{(\mathrm{int},1)}=-2\eta_{r}A_{ij}.$ (87)
@@
$\sigma_{[ij]}^{(\mathrm{int},1)}=\eta_{r}\varepsilon_{ijk}(2\omega_{0k}-\zeta_{k}),\qquad-\varepsilon_{ijk}\sigma_{jk}^{(\mathrm{int},1)}=2\eta_{r}(\zeta_{i}-2\omega_{0i}).$ (88)
@@ V.4 Couple stress at first order
The first-order couple stress is obtained from Eq. (47):
$m_{ij}^{(1)}=\int I\omega_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (97)
Since $\omega_{i}=\Omega_{i}+\omega_{0i}$ and $\int c_{j}f^{(0)}\,\mathrm{d}\Gamma=0$, only $I\Omega_{i}c_{j}$ matters. Introduce the microscopic kernel
@@
$m_{ij}^{(1)}=\int Q_{ij}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (99)
Only the $q$, $b_{i}$, and $E_{ij}$ sectors contribute. Rotational invariance therefore implies the general form
@@
$\lambda_{0}$ $=\frac{1}{3}\int Q_{kk}f^{(0)}Y\,\mathrm{d}\Gamma,$ (101)
$\lambda_{1}$ $=\frac{1}{6}\varepsilon_{ijk}\int Q_{ij}f^{(0)}Z_{k}\,\mathrm{d}\Gamma,$ (102)
$\lambda_{2}$ $=\frac{1}{5}\int\left\langle Q_{ij}\right\rangle f^{(0)}Z_{ij}\,\mathrm{d}\Gamma.$ (103)
@@ V.5 Divergence of the couple stress
The exact spin balance (54) therefore becomes, after the standard micropolar rewriting,
@@ VI.3 Rotational-viscosity sector
The rotational-viscosity sector is subtler. As Eq. (74) showed, the symmetric kinetic stress kernel $mc_{i}c_{j}$ cannot generate an axial stress. The coefficient $\eta_{r}$ therefore belongs to an intrinsic or collisional transfer functional. Denote the linear intrinsic stress functional by $\mathfrak{S}_{ij}[\phi]$. Restricting to the axial channel gives
$\sigma_{[ij]}^{(\mathrm{int},1)}=\mathfrak{S}_{[ij]}[X_{k}a_{k}].$ (119)
@@
where $\mathcal{S}^{(\mathrm{ax})}_{i}$ denotes the net axial source produced by the intrinsic/collisional transfer mechanism and by the retained-spin extended manifold.
@@ VI.4 What is closed and what is not
  • The exact balance laws are fully derived.
@@ VII.2.2 Collisional production of the spin density
The first term is odd under $\bm{g}\mapsto-\bm{g}$ and does not contribute in the homogeneous isotropic reference state. The second term produces the linear relaxation.
The collisional production of the spin density is obtained by multiplying the collision operator by $I\omega_{i}$ and symmetrizing over the two collision partners. The result is
@@ VII.2.3 Angular average
A further isotropic average over the direction of $\bm{g}$ uses $\left\langle\hat{g}_{i}\hat{g}_{j}\right\rangle=\delta_{ij}/3$, giving
$\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right\rangle_{\hat{\bm{g}}}=-\frac{2\pi}{3}\,\left\langle|\bm{g}|\right\rangle\,\delta_{ij}.$ (142)
@@
$\left\langle|\bm{g}|\right\rangle=4\sqrt{\frac{k_{B}T}{\pi m}},$ (143)
@@ VII.2.4 Relaxation rate and constitutive matching
Using Eq. (134) in Eq. (135), dropping the odd term, and using the fact that $\bm{g}$ and $\bm{\Omega}_{+}$ are statistically independent in the reference state, we obtain
@@
$=-\frac{2\pi a^{2}I}{3(K+1)}\left\langle|\bm{g}|\right\rangle\int f_{\mathrm{h}}f_{\mathrm{h},1}\,\Omega_{+,i}\,\mathrm{d}\Gamma\,\mathrm{d}\Gamma_{1}.$ (144)
@@ VII.3.1 Setup and first-order ansatz
$m_{zy}$ $=I\int\Omega_{z}c_{y}f^{(1)}_{\mathrm{sd}}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=IA_{m}g_{s}\int\Omega_{z}^{2}c_{y}^{2}f^{(0)}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=IA_{m}g_{s}n\left\langle\Omega_{z}^{2}\right\rangle\left\langle c_{y}^{2}\right\rangle$
$=IA_{m}g_{s}n\left(\frac{k_{B}T}{I}\right)\left(\frac{k_{B}T}{m}\right)$
@@ VII.3.2 Pair mode and collision kinematics
Define the pair mode
@@
$Y_{ij}=\frac{1}{2}\Omega_{+,i}V_{j}+\frac{1}{2}\Omega_{-,i}g_{j}.$ (160)
@@
$\Delta Y_{ij}=\frac{1}{2}(\Delta\Omega_{+,i})V_{j}+\frac{1}{2}\Omega_{-,i}(\Delta g_{j}).$ (163)
@@ VII.3.3 Relaxation rate as a collision bracket
The linear relaxation rate of the mode is defined by
@@
To evaluate the numerator, introduce
$A_{ij}:=\frac{1}{2}\Omega_{+,i}V_{j},\qquad B_{ij}:=\frac{1}{2}\Omega_{-,i}g_{j},$ (168)
@@
$\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\Delta A_{ij})A_{ij}\right\rangle_{0}$ $=\frac{1}{4}\left\langle V_{j}^{2}\right\rangle_{0}\frac{2}{K+1}\left\langle\Omega_{+,i}\Omega_{+,\ell}\right\rangle_{0}\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{\ell}-\delta_{i\ell})\right\rangle_{0}$
$=\frac{1}{4}(2v_{T}^{2})\frac{2}{K+1}(2\omega_{T}^{2})\left(-\frac{2\pi}{3}\left\langle|\bm{g}|\right\rangle\right)$
@@
In the second line we used Eqs. (142) and (143).
@@ IX Discussion and conclusion
First, the exact balance laws follow directly from the Boltzmann–Curtiss equation once the collision invariants are specified. The intrinsic-spin balance is most transparent when derived from total angular momentum and then reduced by subtracting the orbital part.
Second, when the mean spin $\bm{\omega}_{0}$ is retained explicitly, the relevant first-order Chapman–Enskog calculation is a generalized one in the sense of extended hydrodynamics. In the bookkeeping adopted here, the residual collisional production of the retained-spin quasi-equilibrium manifold is placed at $O(\varepsilon)$, so the retained-spin reference manifold is quasi-equilibrium rather than the strict null manifold of the collision operator.
Third, the irreducible decomposition of the first-order source cleanly separates the shear, bulk, spin-diffusion, and axial mismatch channels. The symmetric kinetic stress contains no axial part. As a result, the rotational viscosity $\eta_{r}$ must belong to an intrinsic or collisional transfer mechanism.

### 13 Apr 2026 v2

@@ IV.3 Linearized first-order equation
The Chapman–Enskog matching conditions keep the retained fields entirely in $f^{(0)}$; in particular,
$\int f^{(1)}\,\mathrm{d}\Gamma=0,\qquad\int c_{i}f^{(1)}\,\mathrm{d}\Gamma=0,\qquad\int\Omega_{i}f^{(1)}\,\mathrm{d}\Gamma=0.$ (67)
@@
where the source $\mathcal{S}$ collects the $O(\varepsilon)$ streaming contributions together with the matched-order residual axial collisional-transfer contribution described above. The standard density, temperature, and heat-flux sectors follow the classical Chapman–Enskog pattern [2, 22]. Here we focus on the sectors relevant to the stress and couple stress.
@@ IV.4 Useful derivatives of the local Maxwellian
The last term $\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}$ is written separately because the axial mismatch sector is not generated by the symmetric one-particle kinetic stress alone. In the matched ordering described above, this symbol also absorbs the $O(\varepsilon)$ residual axial relaxation of the retained-spin quasi-equilibrium manifold. The axial sector therefore belongs to the collisional-transfer channel and will be treated formally in Sec. VI.
@@ V.2 Total stress at first order
The kernel $c_{i}c_{j}f^{(0)}$ is symmetric in $(i,j)$ and even under $\bm{\Omega}\mapsto-\bm{\Omega}$. The axial-mismatch and spin-gradient sectors in Eq. (86) are odd in $\bm{\Omega}$ and therefore integrate to zero against this kernel; in particular, no $q\,\delta_{ij}$ or $E_{ij}$ contribution survives in the kinetic stress. Thus only the symmetric velocity-gradient sectors contribute to $\sigma_{ij}^{(k,1)}$, and rotational invariance forces the result to be of the form
@@
The antisymmetric channel belongs to the collisional-transfer part of the total stress. Introduce the axial mismatch tensor
@@
Isotropy and parity imply that the first-order intrinsic stress in this channel must itself be antisymmetric and proportional to $A_{ij}$, so we define
$\sigma_{ij}^{(\mathrm{int},1)}=-2\eta_{r}A_{ij}.$ (92)
@@
$\sigma_{ij}^{(\mathrm{int},1)}=\eta_{r}\varepsilon_{ijk}(2\omega_{0k}-\zeta_{k}),\qquad-\varepsilon_{ijk}\sigma_{jk}^{(\mathrm{int},1)}=2\eta_{r}(\zeta_{i}-2\omega_{0i}).$ (93)
@@ V.4 Couple stress at first order
The leading one-particle contribution to the first-order couple stress is obtained from Eq. (49):
$m_{ij}^{(k,1)}=-\int I\omega_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (102)
Writing $\omega_{i}=\Omega_{i}+\omega_{0i}$ gives
$m_{ij}^{(k,1)}=-\int I\Omega_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma-\omega_{0i}\int Ic_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (103)
The second term vanishes by the Chapman–Enskog matching condition $\int c_{j}f^{(1)}\,\mathrm{d}\Gamma=\int c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma=0$ from Eq. (67). Introduce the microscopic kernel
@@
$m_{ij}^{(k,1)}=-\int Q_{ij}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (105)
At dilute first order we represent the total couple stress by this leading one-particle contribution, so only the $q$, $b_{i}$, and $E_{ij}$ sectors contribute. Rotational invariance therefore implies the general form
@@
$\lambda_{0}$ $=-\frac{1}{3}\int Q_{kk}f^{(0)}Y\,\mathrm{d}\Gamma,$ (107)
$\lambda_{1}$ $=-\frac{1}{6}\varepsilon_{ijk}\int Q_{ij}f^{(0)}Z_{k}\,\mathrm{d}\Gamma,$ (108)
$\lambda_{2}$ $=-\frac{1}{5}\int\left\langle Q_{ij}\right\rangle f^{(0)}Z_{ij}\,\mathrm{d}\Gamma.$ (109)
@@ V.5 Divergence of the couple stress
The exact spin balance (58) therefore becomes, after inserting the first-order constitutive forms for $m_{ij}$ and $\sigma_{ij}$,
@@ VI.3 Rotational-viscosity sector
The rotational-viscosity sector is subtler. As Eq. (79) showed, the symmetric kinetic stress kernel $mc_{i}c_{j}$ cannot generate an axial stress. The coefficient $\eta_{r}$ therefore belongs to the collisional-transfer part of the total stress. Denote the corresponding linear stress functional by $\mathfrak{S}_{ij}[\phi]$. In the present axial channel this gives
$\sigma_{ij}^{(\mathrm{int},1)}=\mathfrak{S}_{ij}[X_{k}a_{k}].$ (125)
@@
where $\mathcal{S}^{(\mathrm{ax})}_{i}$ denotes the net axial source produced by the collisional-transfer mechanism and by the retained-spin extended manifold.
@@ VI.4 What is closed and what is not
  • The one-particle balance identities are fully derived, and the exact finite-size antisymmetric-stress torque channel is made explicit.
@@ VII.2.2 Collisional production of the spin density
The first term does not contribute after the incoming-hemisphere average over $\bm{k}$: for fixed $\bm{g}$, the integral $\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\bm{k}\times\bm{g})$ vanishes by axial symmetry around $\bm{g}$. The second term produces the linear relaxation.
The dilute collisional production of the spin density may be estimated directly from the pair-collision change in the spin sum and symmetrization over the two collision partners. This gives
@@ VII.2.3 Angular average
For fixed $|\bm{g}|$, an isotropic average over the direction of $\bm{g}$ uses $\left\langle\hat{g}_{i}\hat{g}_{j}\right\rangle_{\hat{\bm{g}}}=\delta_{ij}/3$, giving
$\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right\rangle_{\hat{\bm{g}}}=-\frac{2\pi}{3}|\bm{g}|\,\delta_{ij}.$ (148)
A subsequent Maxwellian average over $|\bm{g}|$ then gives
$\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right\rangle_{0}=-\frac{2\pi}{3}\,\left\langle|\bm{g}|\right\rangle_{0}\,\delta_{ij}.$ (149)
@@
$\left\langle|\bm{g}|\right\rangle_{0}=4\sqrt{\frac{k_{B}T}{\pi m}},$ (150)
@@ VII.2.4 Relaxation rate and constitutive matching
Using Eq. (140) in Eq. (141), discarding the $\bm{k}\times\bm{g}$ term whose incoming-hemisphere average vanishes by axial symmetry around $\bm{g}$, and using the fact that $\bm{g}$ and $\bm{\Omega}_{+}$ are statistically independent in the reference state, we obtain
@@
$=-\frac{2\pi a^{2}I}{3(K+1)}\left\langle|\bm{g}|\right\rangle_{0}\int f_{\mathrm{h}}f_{\mathrm{h},1}\,\Omega_{+,i}\,\mathrm{d}\Gamma\,\mathrm{d}\Gamma_{1}.$ (151)
@@ VII.3.1 Setup and first-order ansatz
$m_{zy}$ $=-I\int\Omega_{z}c_{y}f^{(1)}_{\mathrm{sd}}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=-IA_{m}g_{s}\int\Omega_{z}^{2}c_{y}^{2}f^{(0)}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=-IA_{m}g_{s}n\left\langle\Omega_{z}^{2}\right\rangle\left\langle c_{y}^{2}\right\rangle$
$=-IA_{m}g_{s}n\left(\frac{k_{B}T}{I}\right)\left(\frac{k_{B}T}{m}\right)$
@@
The relaxation calculation below concerns the single transverse component $X_{zy}$. By isotropy, one may equivalently evaluate any fixed Cartesian component pair of the same type, and we may take $(i,j)=(z,y)$ throughout. For notational compactness we nevertheless keep the symbols $i$ and $j$ in the pair-mode algebra, but from this point onward they are fixed component labels rather than dummy indices: Einstein summation is not being used for $i$ or $j$ (that is, there is no sum over $i$ or $j$). Only explicitly repeated auxiliary indices such as $\ell$ are summed.
@@ VII.3.2 Pair mode and collision kinematics
Define the pair mode for such a fixed component pair
@@
$Y_{ij}=\frac{1}{2}\Omega_{+,i}V_{j}-\frac{1}{2}\Omega_{-,i}g_{j}.$ (167)
@@
$\Delta Y_{ij}=\frac{1}{2}(\Delta\Omega_{+,i})V_{j}-\frac{1}{2}\Omega_{-,i}(\Delta g_{j}).$ (170)
@@ VII.3.3 Relaxation rate as a collision bracket
The linear relaxation rate of this fixed-component mode is defined by
@@
To evaluate the numerator for this fixed component pair, introduce
$A_{ij}:=\frac{1}{2}\Omega_{+,i}V_{j},\qquad B_{ij}:=-\frac{1}{2}\Omega_{-,i}g_{j},$ (175)
@@
$\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\Delta A_{ij})A_{ij}\right\rangle_{0}$ $=\frac{1}{4}\left\langle V_{j}^{2}\right\rangle_{0}\frac{2}{K+1}\sum_{\ell=1}^{3}\left\langle\Omega_{+,i}\Omega_{+,\ell}\right\rangle_{0}\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{\ell}-\delta_{i\ell})\right\rangle_{0}$
$=\frac{1}{4}(2v_{T}^{2})\frac{2}{K+1}(2\omega_{T}^{2})\left(-\frac{2\pi}{3}\left\langle|\bm{g}|\right\rangle_{0}\right)$
@@
In the second line we used the Maxwellian-averaged tensor identity (149) together with Eq. (150).
@@ IX Discussion and conclusion
First, the one-particle balance identities follow directly from the Boltzmann–Curtiss equation once the collision invariants are specified, while for finite-size rough particles the exact local intrinsic-spin balance must keep the antisymmetric-stress torque explicit. The intrinsic-spin balance is most transparent when derived from total angular momentum and then reduced by subtracting the orbital part.
Second, when the mean spin $\bm{\omega}_{0}$ is retained explicitly, the relevant first-order Chapman–Enskog calculation is a generalized one in the sense of extended hydrodynamics. In the bookkeeping adopted here, the residual axial relaxation of the retained-spin manifold is placed at $O(\varepsilon)$ within the collisional-transfer channel, so the retained-spin reference manifold is quasi-equilibrium rather than a strictly instantaneous elimination of spin.
Third, the irreducible decomposition of the first-order source cleanly separates the shear, bulk, spin-diffusion, and axial mismatch channels. The symmetric one-particle kinetic stress contains no axial part, and the spin-gradient sectors are excluded from it by parity. As a result, the rotational viscosity $\eta_{r}$ belongs to the collisional-transfer stress torque channel.